OpenStudy (theviper):

The Quadratic Polynomial whose zeroes are \(\Large{\frac{3}{5}}\) & \(\Large{\frac{-1}{2}}\) is ???? @AravindG

OpenStudy (aravindg):

(x-a)(x-b)=0 Has its zeros a and b

OpenStudy (aravindg):

Use this^

OpenStudy (theviper):

I got \[\Large{\frac{10x^2-x-3}{10}}\] but book says \[\Large{10x^2-x-3}\]

OpenStudy (aravindg):

(x-3/2)(x+1/2)=0 (5x-3)(2x+1)=0 10x^2-x-3=0 Yes your book is right!

OpenStudy (aravindg):

Infact you are right too...Multiply by 10 on both sides

OpenStudy (theviper):

Oh 10 is multiplied by '0' on the RHS Thanks Bhaiya :)

OpenStudy (aravindg):

yw :)